# Concentration Inequalities for Suprema of Empirical Processes with Dependent Data via Generic Chaining with Applications to Statistical Learning

### v1

@@ A.2 Proofs for Section 4
$\mathbb{P}(Y_{i}\geq\varepsilon)=\mathbb{P}(|\widetilde{X}_{i}|\geq\varepsilon+\zeta)\leq 2\exp(-(\varepsilon+\zeta)^{\alpha})$
$\quad\leq 2\exp(-\varepsilon^{\alpha}-\zeta^{\alpha})=\exp(-\varepsilon^{\alpha})~.$
@@
Since, $X_{\bm{\theta}}$ is a separable process there exists a countable dense subset $\tilde{\Theta}\subset\Theta$ such that
@@
Proof of Proprosition 4.4.
The proof is based on a straightforward generalization of the arguments in (Talagrand, 2005, pp. 22–23). For the metric space $(\Theta,d_{\theta})$ we define the $k$-th entropy number as $e_{k}(\Theta)=\inf_{\mathcal{A}_{k}}\sup_{\bm{\theta}\in\Theta}\Delta(A_{k}(\bm{\theta}))$. That is the smallest radius such that $\Theta$ can be covered by at most $2^{2^{k}}$ balls of that radius. Equivalently,
@@ A.3 Properties of sub-Weibull Random Variables
When $\alpha\geq 1$, the sub-Weibull quasi-norm is in fact a norm, and the result follows directly from the triangle inequality. Vershynin (2026) shows that the sub-Weibull quasi-norm is a true norm for $\alpha=1$ and $\alpha=2$, corresponding to sub-exponential and sub-Gaussian random variables ∎
@@
$\log\left(\mathbb{E}\left[\left|1+\frac{sX_{i}}{4}\right|^{p}\right]\right)\leq\begin{cases}N_{\alpha}^{*}(p|s|),&\text{if }p|s|\geq 2\\ p^{2}s^{2},&\text{if }p|s|<2\end{cases}~,$
@@
Let us now focus on the case $\alpha=1$. It is straightforward to verify that

### v2

@@ A.2 Proofs for Section 4
$\mathbb{P}(Y_{i}\geq\varepsilon)=\mathbb{P}(|\widetilde{X}_{i}|\geq\varepsilon+\zeta)\leq 2\exp(-(\varepsilon+\zeta)^{\alpha})\leq 2\exp(-\varepsilon^{\alpha}-\zeta^{\alpha})=\exp(-\varepsilon^{\alpha})~,$
@@
Since $X_{\bm{\theta}}$ is a separable process there exists a countable dense subset $\tilde{\Theta}\subset\Theta$ such that
@@
$r$
Proof of Proposition 4.4.
The proof is based on a straightforward generalization of the arguments in (Talagrand, 2005, pp. 22–23). For the metric space $(\Theta,d_{\theta})$ we define the $k$-th entropy number as $e_{k}(\Theta)=\inf_{\mathcal{A}_{k}}\sup_{\bm{\theta}\in\Theta}\Delta(A_{k}(\bm{\theta}))$. It is the smallest radius such that $\Theta$ can be covered by at most $2^{2^{k}}$ balls of that radius. Equivalently,
@@ A.3 Properties of sub-Weibull Random Variables
When $\alpha\geq 1$, the sub-Weibull quasi-norm is in fact a norm, and the result follows directly from the triangle inequality. Vershynin (2026) shows that the sub-Weibull quasi-norm is a true norm for $\alpha=1$ and $\alpha=2$, corresponding to sub-exponential and sub-Gaussian random variables. ∎
@@
$\log\left(\mathbb{E}\left[\left|1+\frac{sX_{i}}{4}\right|^{p}\right]\right)\leq\begin{cases}N_{\alpha}^{*}(p|s|),&\text{if }p|s|\geq 2\\ p^{2}s^{2},&\text{if }p|s|<2\end{cases}$~
@@
Let us focus on the case $\alpha=1$. It is straightforward to verify that